OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -1/2x + 5. A)y = -2x + 2 B)y = 2x - 14 C)y = 1/2x - 8 D)y = -1/2x - 4

OpenStudy (zzr0ck3r):

we know it will have same slope of -1/2 because its parallel so we know it will be of the form y=-(1/2)x+b so use the given point to find b plug -6 for y and 4 for x -6=-(1/2)4+b -6=-2+b -4=b so y=-(1/2)X-4

OpenStudy (anonymous):

Thanks i knew a and b were wrong